OpenStudy (anonymous):

Find the exact value of cot 15degrees. (Use Half Angle Formula)

OpenStudy (anonymous):

An easier alternative would be to use the subtraction formula for \[\cos \left( 45-30 \right)^{o}\div \sin \left( 45-30 \right)^{o}\]

OpenStudy (anonymous):

Gotcha. Thanks for your help

OpenStudy (anonymous):

you're welcome.

OpenStudy (y2o2):

\[2\sin²(θ) =1-\cos(2θ)\] \[2\cos²(θ) =1+\cos(2θ)\] \[\cot²(15) = \] \[ (1+ \cos30) \over (1- \cos30)\]